some guy rowed 18 miles down the delaware river in two hours but the return trip took him four and a half hours find the rate this guy can row in still water and find the rate of the current
Let X = rate guy can row in still water
Let Y = rate of current

D = r . T

Down stream x plus y
upstream x - y

Answer by Alan3354(69443)

Downstream speed = 18/2 = 9 mph
Upstream = 18/4.5 = 4 mph
--------
Rowing speed = the average = (9+4)/2 = 6.5 mph
Current = the difference = 9 - 6.5 = 2.5 mph